OpenStudy (micahm):

Consider the numbers 4/13and 3/10 . Find a fraction that is between these two numbers. Show all work. can i have step by step help

OpenStudy (micahm):

this sign / is a fraction

OpenStudy (solomonzelman):

add them and divide by 2.

OpenStudy (solomonzelman):

(just a simple average problem)

OpenStudy (micahm):

what numbers

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE 4 }{\LARGE13 } }\) and \(\large\color{black}{ \frac{\LARGE 3}{\LARGE10 }}\) would have an average of: \(\large\color{black}{ \left(\begin{matrix} \frac{\LARGE 4 }{\LARGE13 } +\frac{\LARGE 3}{\LARGE10 } \\ \end{matrix}\right)\div2 }\)

OpenStudy (micahm):

7 23

OpenStudy (solomonzelman):

( just like any \(\large\color{black}{ \rm B }\) and \(\large\color{black}{ \rm C }\) have an average of: \(\large\color{black}{ ({\rm B }+{\rm C }) \div 2}\) )

OpenStudy (solomonzelman):

incorrectly added the fractions

OpenStudy (solomonzelman):

you have to find the common denominator, between 13 and 10. Can you tell me what is it?

OpenStudy (micahm):

like what number they both have the you can divid

OpenStudy (solomonzelman):

like the smallest number that you can divide by 10 and by 13 (without any remainder).

OpenStudy (micahm):

2 or 3

OpenStudy (micahm):

@Secret-Ninja

OpenStudy (micahm):

@paki

OpenStudy (micahm):

@Compassionate